OpenStudy (juliannegirltomboy):

6th GRADE MATH :Ratios and Ratio Language

OpenStudy (sagewilson):

Post your question please.

OpenStudy (juliannegirltomboy):

ok one moment

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

H + T = ?

OpenStudy (juliannegirltomboy):

=75

OpenStudy (skullpatrol):

Nope.

OpenStudy (sagewilson):

Okay. he slipped it 45 times, since he got heads 30 times this would give you 30/45 meaning he got tails 15 times. since tails is smaller than the amount of heads. what would your ratio look like?

OpenStudy (sagewilson):

*flipped

OpenStudy (juliannegirltomboy):

a?

OpenStudy (skullpatrol):

H + T = 45

OpenStudy (sagewilson):

I'm not asking you to answer it, I'm just asking you to tell me what you think the ratio would look like unsimplified.

rebeccaxhawaii (rebeccaxhawaii):

how much is 30 outta 45

OpenStudy (juliannegirltomboy):

15

OpenStudy (skullpatrol):

H = 30

OpenStudy (skullpatrol):

T = ?

OpenStudy (juliannegirltomboy):

15

OpenStudy (skullpatrol):

H/T = ?

OpenStudy (juliannegirltomboy):

2

OpenStudy (skullpatrol):

As a ratio 2 = 2/1

OpenStudy (juliannegirltomboy):

thanks

OpenStudy (juliannegirltomboy):

can you help with some more

OpenStudy (skullpatrol):

Post again please

OpenStudy (juliannegirltomboy):

ok

OpenStudy (juliannegirltomboy):

is it A.

OpenStudy (skullpatrol):

Yes

OpenStudy (mysticfury):

you are correct @Juliannegirltomboy

OpenStudy (juliannegirltomboy):

i have 6 more

OpenStudy (juliannegirltomboy):

c?

OpenStudy (skullpatrol):

Yes

OpenStudy (juliannegirltomboy):

24TO8?

OpenStudy (skullpatrol):

Yes